Assume the current U.S. dollar-British spot rate is $1.3063=£. If the current nominal one-year interest rate in the U.S. is 1.5% and the comparable rate in Britain is 0.75%, what is the approximate forward exchange rate for 360 days?

A. £1.2965/$
B. £0.7598/$
C. $1.2965/£
D. £1.3161/$
Answer #1

For ward Exchange rate for 360 days =Spot Rate*(1+US Rate)/(1+Pound Rate) =1.3063*(1+1.5%)/(1+0.75%) =1.3160243

Forward Exchange Rate pound to dollar =1/1.3160243 =0.7598 Pound/Dollar